Hello! I need help on this question. I understand all the speed x time equals distance and so, but i don't quite get this kinds of questions. A man walks to work at 3 1/2 kmph and is one minute late. If he walked at 3 1/3 kmph he would have been three minutes late. What is the distance to his office? Thanks :)
Answer by mananth(16949)

let the right time be t
speed = 3 1/2 kmph time taken = t+(1/60) hours
speed =3 1/3 kmph time taken = t + (3/60) hours
in both cases distance is the same
d=rt
3 1/2 *( t+(1/60) )= distance
3 1/3 (t + (3/60))= distance


3 1/2 *( t+(1/60) )= 3 1/3 (t + (3/60))

7/2((60t+1)/60) = 10/3((60t+3)/60)

cross multiply
21(60t+1)/60 = 20(60t+3)/60
21(60t+1)=20(60t+3)
1260t+21=1200t+60
60t=60-21
60t=39
t= 39/60
t= 39 minutes
With speed of 3 1/2 he takes one minute more => 39+1=40 minutes
d=rt
d= (7/2) *(40/60)
d= 7/3 =2.33 km